# When Does Divide and Conquer Work for Long Context LLM? A Noise Decomposition Framework

### v1

@@ 4 A simple implementation of the framework
Planner: The planner orchestrates how the input is divided, what instructions each worker agent receives, and how the manager agent is prompted to unify the partial outputs. Rather than having a human specify how to break down the input, we direct the planner to do so by following steps:
  1. 1.
    Job Assignment. The planner decides how many chunks to create and which segments each worker agent will process.
  2. 2.
    Prompt Preparation. Based on the task details, the planner modifies the worker prompts so that each worker’s output can be correctly integrated downstream. It also sets up the manager prompt to combine partial results properly.
  3. 3.
    Iterative Refinement. The planner can run a brief evaluation step using some validation data to identify mispredicted cases. It then revises the prompt structure or chunking strategy to reduce errors. If repeated excessively on the same data, there is a risk of overfitting, so the planner typically does only a few refinements.
In this manner, the planner automates much of the chunking and coordination that would otherwise require manual trial and error. As we will see in the experiments, careful planner design can substantially reduce aggregator noise by ensuring each worker agent’s output is in a format that the manager agent can interpret reliably.
@@ 5.1 Settings
We present experiments to assess how the three primary noise components—model noise, task noise, and aggregator noise—affect system performance under different context lengths and across multiple tasks. Section 3.4 indicates that $\phi_{\mathrm{model}}$ grows more than linearly with input length, and that synergy across subproblems ($\Phi_{\mathrm{task}}$) can impose strong requirements on any mechanism for merging partial outputs. Our experiments test the ideas on six tasks with different data scales and use several agents.
@@
We experiment on six diverse tasks including: Key-Value Retrieval, Math Find Number, Summarization, Dialogue Character Inference, and Open Question QA with and without choices. There tasks are based on InfiniteBench Zhang et al. (2024a) and LongBench-V2 Bai et al. (2024) but we have modified the generation and prepared different lengths of these tasks.
  • Key-Value Retrieval (KV) We randomly generated Key-Value pairs. For a given key, the task is to retrieve the value associated. This task only evaluates the capability of retrieval. The evaluation metric is accuracy. We prepare synthetic KV task to have lengths ranging from 1K to 128K.
  • Math Find Number (Math) We randomly generated a long list of integers following Gaussian distribution and the task is to find the Kth largest or smallest number in this list. Each query is slightly different in K. This task evaluates the memory and mathematical reasoning. The evaluation metric is accuracy. We prepare synthetic Math task to have lengths ranging from 1K to 128K.
  • Summarization (Sum) We provide English texts and the task is to summarize them. The ground truth summarization comes from the InfiniteBench. This task evaluates the langauge summarization. The evaluation metric is ROUGE score.
  • Open Question QA (QA-IB and QA-LB) We provide English text and some open questions for question answering. These questions are very diverse and the task requires LLM to do multi-hop reasoning to answers. The task QA-IB comes from InfiniteBench and doesn’t come with choices and the task QA-LB comes from LongBench-V2 and is provided with four choices. The evaluation metrics are respectively F1 score and accuracy.
  • Dialogue Character Inference (Char) We provide English dialogue scripts between many characters. One of the characters is masked and the task is to infer the name of the masked person. This task evaluates the capacity of language reasoning and will require LLM to focus on the interaction of people discussions to understand the relation. The evaluation metric is accuracy. This task comes from InfiniteBench.
@@
We have experimented with diverse agents choices from commercial OpenAI models and open sourced models from Meta and Alibaba. gpt-4o-2024-08-06 (gpt4o) 128K model from OpenAI. gpt-4o-mini-2024-07-18 (gpt4omini) 128K model from OpenAI. Llama-3.1-70B-Instruct (llama70b) 128K 70B LLM model from Meta Grattafiori (2024). Llama-3.2-3B-Instruct (llama3b) 128K 3B LLM model from Meta Grattafiori (2024). Qwen2.5-72B-Instruct (qwen72b) 32K 72B LLM model from Alibaba QWen Yang et al. (2024).
In the experiments where divide and conquer agents are implemented, the manager and worker agents are homogeneous and the planner agent is QWen72b. The temperature is set to 0 to minimize stochasticity during decoding.
@@
In the following sections, we discuss respectively model noise, task noise and aggregator noise. The robustness of these observations and the practical applicability of our framework are further underscored by comprehensive supplementary analyses (detailed in Appendices D-J), which include utility of the framework, explorations of chunking variants, comparisons with RAG, evaluations across diverse model architectures, and detailed data for our main findings.
5.2 Model noise
@@
We estimate the single-agent model noise for input $x$ as $\phi_{\mathrm{model}}(x)\;\approx\;\mathcal{D}\bigl{(}f_{\mathrm{single}}(x),\,f^{*}(x)\bigr{)},$ where $f_{\mathrm{single}}(x)$ is the output of a single model given the entire input, and $\mathcal{D}(\cdot,\cdot)$ is a distance metric such as cross-entropy or $|1-\text{accuracy}|$. By evaluating this for different input lengths $T=\|x\|$, we see how the model’s performance declines as context grows. In Figures 2(b) (Math) and 2(a) (KV), each agent processes the entire input directly (up to 128K tokens). We measure accuracy at each length; the detailed numerical scores supporting these figures are provided in Appendix D.
(a) Single agent model noise w.r.t input length: KV Task
@@
(b) Single agent model noise w.r.t input length: Math Task
@@
For KV retrieval (Figure 2(a)), most agents show a downward trend in accuracy as input length increases, consistent with Assumption 3.1, which posits that noise grows more than linearly. Once the length exceeds tens of thousands of tokens, performance rapidly decays. Similar observations are seen for Math (Figure 2(b)), where the loss of accuracy at longer contexts is stark for gpt4omini and llama70b. Interestingly, llama3b is already too weak at shorter lengths, so its performance saturates at near-random levels. These results reinforce our theoretical statement that $\phi_{\mathrm{model}}(x)$ can become the dominant error term for long $x$. For tasks that do not require intense synergy across the entire input, splitting the input into shorter segments (i.e., reducing $L$ per worker) should mitigate this growth, an idea formalized in Thm 3.2.
5.3 Task noise
Next, we examine how multi-agent performance depends on $\Phi_{\mathrm{task}}$ and $\Phi_{\mathrm{model}}$. We use the approximate measures:
$\Phi_{\mathrm{model}}\;\approx\;\mathcal{D}\bigl{[}f_{\mathrm{D\&C}}(x;\,\mathrm{optimal\text{-}n}),\,f_{\mathrm{D\&C}}(x;\,\mathrm{bad\text{-}n})\bigr{]},$
$\Phi_{\mathrm{task}}\;\approx\;\int_{2}^{n}\mathcal{D}\bigl{[}f_{\mathrm{D\&C}}(x;\lfloor i\rfloor),\,f_{\mathrm{single}}(x)\bigr{]}\;\mathrm{d}i.$
Here, $\Phi_{\mathrm{model}}$ is captured by the gap between choosing an effective chunk size versus a poor one, while $\Phi_{\mathrm{task}}$ is reflected by how the system output diverges from the single-agent baseline as we increase the number of chunks from 2 to $n$.
@@ 5.3 Task noise
Figure 3: Joint effect of task noise $\Phi_{\mathrm{task}}$ and model noise $\Phi_{\mathrm{model}}$. According to the discussion in Sec 3.4, (a) has negligible $\Phi_{\mathrm{model}}$ and $\Phi_{\mathrm{task}}$;(b)-(e) has dominating model noises $\Phi_{\mathrm{model}}$; (f) has dominating task noise $\Phi_{\mathrm{task}}$.
Figure 3 illustrates the performance of five agents on the 128K version of each task, varying the chunk size from 1K to 64K tokens. The comprehensive performance metrics for all tasks and models underpinning this figure are available in Appendix E. We observe three patterns that match the three regimes described in Section 3.4:
  1. 1.
    Low task noise (KV). When cross-chunk synergy is minimal, splitting yields similar results regardless of $n$. This matches the “low task noise, low model noise” regime, where $\Phi_{\mathrm{task}}$ is nearly zero and $\phi_{\mathrm{model}}$ remains small, so aggregator choices have little impact.
  2. 2.
@@
    Dominating model noise (Math, QA, Sum). In these tasks, large input length makes single-shot usage prone to confusion, so $\phi_{\mathrm{model}}$ becomes significant. However, $\Phi_{\mathrm{task}}$ is moderate enough that a basic aggregator can handle partial results effectively. Splitting into smaller chunks thus reduces per-chunk confusion and outperforms a single-shot approach. This behavior aligns with Thm 3.2, showing that superlinear growth of $\phi_{\mathrm{model}}$ can be contained with divide-and-conquer.
  3. 3.
    Dominating task noise (Char). In this scenario, cross-chunk interactions are so extensive that partial outputs cannot capture the global context unless the aggregator reintroduces nearly the entire input. If the aggregator does not do so, performance remains low. This corresponds to the “dominating task noise” regime, where $\Phi_{\mathrm{task}}$ is large enough to overshadow the benefits of splitting.
@@
We also explored whether introducing a small token overlap between chunks could mitigate task noise effects. However, as detailed in Appendix G, a 1K token overlap yielded mixed and generally marginal benefits for the Llama-70B model on 128K tasks, suggesting it does not fundamentally alter these noise trade-offs for the scenarios tested. Furthermore, to assess the generalizability of our noise framework, we evaluated additional diverse models across various tasks, as detailed in Appendix J. These results confirm that while effective context length (e.g., per Ruler benchmark) influences performance, the observed noise patterns and the utility of our decomposition are consistent, with task complexity significantly interacting with these factors across different architectures.
@@
Overall, these empirical data show that the relative magnitudes of $\phi_{\mathrm{model}}$ and $\Phi_{\mathrm{task}}$ determine whether chunking helps or hurts. Moderate-synergy tasks benefit most from a chunk-based approach, validating our theoretical insights in Section 3.4. Apart from the high-synergy character-inference problem, divide-and-conquer methods achieve performance on par with or exceeding the strongest single-shot model, reinforcing the idea that chunking is advantageous when model noise dominates and task noise remains modest.
5.4 Aggregator noise
Finally, we examine $\Phi_{\mathrm{agg}}$, which measures how well the partial results are merged. We estimate by
$\Phi_{\mathrm{agg}}\;\approx\;\int_{2}^{n}\mathcal{D}\bigl{[}f^{\mathrm{optimal\text{-}agg}}_{\mathrm{D\&C}}(x;\lfloor i\rfloor),f^{\mathrm{bad\text{-}agg}}_{\mathrm{D\&C}}(x;\lfloor i\rfloor)\bigl{]}di,$
@@ 5.4 Aggregator noise
Here, we compare a stronger aggregator prompt (planner-based) to a weaker one (manual). Figure 4 contrasts two approaches on the Math and QA-LB tasks for large $T$: (1) a manual aggregator prompt, which simply asks each worker to solve its subproblem in isolation, and (2) a planner-based aggregator prompt, which enforces a structured approach to partial outputs. This yields a visible performance gap (shaded area in Figure 4), showing that aggregator noise can be substantially reduced via more coordinated prompts. The same effect is seen in QA-LB, where the aggregator must reconcile partial answers from each worker.
@@
Figure 4: Aggregator noise across different tasks and models.
These empirical results consolidate Section 3.4 and the aggregator noise assumption (A): a robust aggregator prompt can keep $\Phi_{\mathrm{agg}}$ modest, while naive prompts inflate the final error. The planner effectively refines aggregator instructions to align worker outputs with the manager’s requirements, illustrating how a minimal overhead aggregator can preserve the theoretical advantages of chunking large inputs.

### v2

@@ 4 A simple implementation of the framework
Planner: The planner orchestrates how the input is divided, what instructions each worker agent receives, and how the manager agent is prompted to unify the partial outputs. Rather than having a human specify how to break down the input, we direct the planner to do so by following steps: (1) Job Assignment. The planner decides how many chunks to create and which segments each worker agent will process. (2) Prompt Preparation. Based on the task details, the planner modifies the worker prompts so that each worker’s output can be correctly integrated downstream. It also sets up the manager prompt to combine partial results properly. (3) Iterative Refinement. The planner can run a brief evaluation step using some validation data to identify mispredicted cases. It then revises the prompt structure or chunking strategy to reduce errors. If repeated excessively on the same data, there is a risk of overfitting, so the planner typically does only a few refinements.
Fast chunk-size estimation via sparse sampling
Selecting an appropriate chunk size is critical when model noise dominates. Motivated by our theoretical framework, we adopt a minimal-budget procedure to pick an approximately optimal chunk size without an exhaustive grid search.
Inputs. Candidate chunk sizes $\mathcal{C}$, a small per-configuration sample budget $m$, a development set $\mathcal{D}$ of tasks, and a task-specific metric $\mathsf{M}$.
Procedure. For each $c\in\mathcal{C}$, draw $m$ random documents $S_{c}\subset\mathcal{D}$ (without replacement). Run the D&C pipeline with chunk size $c$ on $S_{c}$ and record $\hat{s}(c)=\frac{1}{m}\sum_{x\in S_{c}}\mathsf{M}(\text{D\&C}(x;c))$. Select $c^{\star}=\arg\max_{c\in\mathcal{C}}\hat{s}(c)$. Deploy the D&C pipeline on the full set using $c^{\star}$.
Complexity and rationale. This reduces the search cost from $O(|\mathcal{D}|\cdot|\mathcal{C}|)$ evaluations to $O(m\cdot|\mathcal{C}|)$ with $m\ll|\mathcal{D}|$. When $\mathcal{L}_{\mathrm{model}}$ dominates and the underlying length-induced degradation $g(L)$ grows superlinearly and is near-monotone in $L$ (Sec. 3.6), the D&C error versus chunk size typically exhibits a clear near-convex optimal region; a few random samples per configuration suffice to localize this optimum.
@@ 5.1 Settings
In our empirical analysis and general discussions, we use the terms task noise, model noise, and aggregator noise as intuitive, macroscopic shorthands for the performance degradations driven by their corresponding formal terms $\mathcal{L}_{\mathrm{task}},\mathcal{L}_{\mathrm{model}},\mathcal{L}_{\mathrm{agg}}$. We present experiments to assess how the three primary noise components—model noise, task noise, and aggregator noise—affect system performance under different context lengths and across multiple tasks. Section 3.6 indicates that the model term $\mathcal{L}_{\mathrm{model}}$ can grow more than linearly with input length, and that cross-chunk dependencies (captured by $\mathcal{L}_{\mathrm{task}}$) can impose strong requirements on any mechanism for merging partial outputs. Our experiments test the ideas on six tasks with different data scales and use several agents.
@@
We experiment on six diverse tasks including: Key-Value Retrieval, Math Find Number, Summarization, Dialogue Character Inference, and Open Question QA with and without choices. These tasks are based on InfiniteBench Zhang et al. (2024b) and LongBench-V2 Bai et al. (2024) but we have modified the generation and prepared different lengths of these tasks. These tasks include: Key-Value Retrieval (KV), Math Find Number (Math), Summarization (Sum) , Open Question QA (QA-IB and QA-LB), Dialogue Character Inference (Char). Detailed task descriptions are provided in Appendix D.
@@
We have experimented with diverse agents choices from commercial OpenAI models and open sourced models from Meta and Alibaba. gpt-4o-2024-08-06 (gpt4o) 128K model from OpenAI. gpt-4o-mini-2024-07-18 (gpt4omini) 128K model from OpenAI. Llama-3.1-70B-Instruct (llama70b) 128K 70B LLM model, Llama-3.2-3B-Instruct (llama3b) 128K 3B LLM model from Meta Grattafiori (2024). Qwen2.5-72B-Instruct (qwen72b) 32K LLM model from Alibaba QWen Yang et al. (2024).
@@
In the experiments where divide and conquer agents are implemented, the manager and worker agents are homogeneous and the planner agent is QWen72b. The temperature is set to 0 to minimize stochasticity during decoding. In the following sections, we study these effects through three empirical lenses: single-agent length-induced degradation, task-term/decomposability effects, and aggregator effects. The robustness of these observations and the practical applicability of our framework are further underscored by comprehensive supplementary analyses (detailed in Appendices F-O), which include utility of the framework, explorations of chunking variants, comparisons with RAG, evaluations across diverse model architectures, and etc.
5.2 Length-Induced Model Degradation
@@
We estimate the single-agent model error for input $x$ directly via the empirical task score: $\mathcal{E}_{\mathrm{single}}(x)=1-S(f_{\mathrm{single}}(x)),$ where $S(\cdot)$ is the task-specific normalized score (e.g., Accuracy). By evaluating this for different input lengths $T=\|x\|$, we see how the model’s performance declines as context grows. In Figures 2(b) (Math) and 2(a) (KV), each agent processes the entire input directly (up to 128K tokens). We measure accuracy at each length; the detailed numerical scores supporting these figures are provided in Appendix F.
(a) Single agent model performance w.r.t input length: KV Task
@@
(b) Single agent model performance w.r.t input length: Math Task
@@
For KV retrieval (Figure 2(a)), most agents show a downward trend in accuracy as input length increases, consistent with the Super-Linear Collapse behavior assumed in Proposition 3.1 (loss growing faster than linearly with context length). Once the length exceeds tens of thousands of tokens, performance rapidly decays. Similar observations are seen for Math (Figure 2(b)), where the loss of accuracy at longer contexts is stark for gpt4omini and llama70b. Interestingly, llama3b is already too weak at shorter lengths, so its performance saturates at near-random levels. These results reinforce our theoretical statement that length-induced degradation can become the dominant error term for long $x$ (the “Brain Fog” regime in Sec. 3.6). For tasks that do not require intense synergy across the entire input, splitting the input into shorter segments (i.e., reducing $L$ per worker) should mitigate this growth, an idea formalized in Proposition 3.1.
5.3 Task Term and Decomposability
Next, we examine how multi-agent performance depends on the task term (cross-chunk dependency / decomposability, $\mathcal{L}_{\mathrm{task}}$) and the model term (length-induced degradation, $\mathcal{L}_{\mathrm{model}}$). Since these decomposition terms are not directly observable from task metrics, we use proxies: we vary the chunk size to control per-worker context length (capturing sensitivity to length-induced model degradation), and we summarize cross-chunk dependency using a proxy based on how D&C outputs deviate from the single-agent baseline over the discrete set of chunk configurations $\mathcal{C}$.
@@ 5.3 Task Term and Decomposability
Figure 3: Joint effect of the task term (decomposability / cross-chunk dependency, $\mathcal{L}_{\mathrm{task}}$) and the model term (length-induced degradation, $\mathcal{L}_{\mathrm{model}}$). According to the discussion in Sec 3.6, (a) has negligible task/model terms; (b)-(e) are model-term dominated; (f) is task-term dominated.
@@
Figure 3 illustrates the performance of five agents on the 128K version of each task, varying the chunk size from 1K to 64K tokens. The comprehensive performance metrics for all tasks and models underpinning this figure are available in Appendix G. We observe three patterns that match the three regimes described in Section 3.6: (1) Low task noise (KV). When cross‐chunk synergy is minimal, splitting yields similar results regardless of $n$. This matches the “low task noise, low model noise” regime, where $\mathcal{L}_{\mathrm{task}}$ is negligible and $\mathcal{L}_{\mathrm{model}}$ remains small, so aggregator choices have little impact. (2) Dominating model noise (Math, QA, Sum). In these tasks, large input length makes single‐shot usage prone to confusion, so $\mathcal{L}_{\mathrm{model}}$ becomes significant. However, $\mathcal{L}_{\mathrm{task}}$ is moderate enough that a basic aggregator can handle partial results effectively. Splitting into smaller chunks thus reduces per‐chunk confusion and outperforms a single‐shot approach. This behavior aligns with Proposition 3.1, showing that the accelerated growth of model error can be contained with divide‐and‐conquer. (3) Dominating task noise (Char). In this scenario, cross‐chunk interactions are so extensive that partial outputs cannot capture the global context unless the aggregator reintroduces nearly the entire input. If the aggregator does not do so, performance remains low. This corresponds to the “dominating task noise” regime, where $\mathcal{L}_{\mathrm{task}}$ is large enough to overshadow the benefits of splitting.
@@
We also explored whether introducing a small token overlap between chunks could mitigate task noise effects. However, as detailed in Appendix I, a 1K token overlap yielded mixed and generally marginal benefits for the Llama-70B model on 128K tasks, suggesting it does not fundamentally alter these noise trade-offs for the scenarios tested. Furthermore, to assess the generalizability of our noise framework, we evaluated additional diverse models across various tasks, as detailed in Appendix L. These results confirm that while effective context length (e.g., per Ruler benchmark) influences performance, the observed noise patterns and the utility of our decomposition are consistent, with task complexity significantly interacting with these factors across different architectures.
@@
Overall, these empirical data show that the relative magnitudes of $\mathcal{L}_{\mathrm{model}}$ and $\mathcal{L}_{\mathrm{task}}$ determine whether chunking helps or hurts. Moderate-synergy tasks benefit most from a chunk-based approach, validating our theoretical insights in Section 3.6. Apart from the high-synergy character-inference problem, divide-and-conquer methods achieve performance on par with or exceeding the strongest single-shot model, reinforcing the idea that chunking is advantageous when the model term dominates and the task term remains modest.
5.4 Aggregator Effects
@@
We examine the aggregator term $\mathcal{L}_{\mathrm{agg}}$, which reflects how well the partial results are merged. Here, we compare a stronger aggregator prompt (planner-based) to a weaker one (manual) over the discrete set of chunk configurations $\mathcal{C}$. Figure 4 contrasts two approaches on the Math and QA-LB tasks for large $T$: (1) a manual aggregator prompt, which simply asks each worker to solve its subproblem in isolation, and (2) a planner-based aggregator prompt, which enforces a structured approach to partial outputs. This yields a visible performance gap (shaded area in Figure 4), showing that aggregation errors can be substantially reduced via more coordinated prompts. The same effect is seen in QA-LB, where the aggregator must reconcile partial answers from each worker. These empirical results consolidate Section 3.6 and our claim that aggregator prompting plays a major role: a robust aggregator prompt can keep $\mathcal{L}_{\mathrm{agg}}$ modest, while naive prompts inflate the final error. The planner effectively refines aggregator instructions to align worker outputs with the manager’s requirements, illustrating how a minimal overhead aggregator can preserve the theoretical advantages of chunking large inputs.
@@
Figure 4: Aggregator errors across different tasks and models.
5.5 Fast estimation of the optimal chunk size
Table 1: Predictive Utility: Optimal Chunk Size Estimation on QA-IB and Summarization (Score (Optimal Chunk Size Found)). Total task length 128K tokens.
Model 3-sample 5-sample 10-sample Optimal after Exhaustive Search
QA-IB
gpt4omini 0.38 (64K) 0.42 (32K) 0.42 (32K) 0.42 (32K)
llama70b 0.55 (2K) 0.63 (16K) 0.63 (16K) 0.63 (16K)
qwen72b 0.40 (2K) 0.48 (16K) 0.48 (8K) 0.48 (8K & 16K)
Sum
gpt4omini 0.15 (16K) 0.14 (8K) 0.14 (8K) 0.15 (4K & 16K)
llama70b 0.23 (32K) 0.24 (16K) 0.28 (8K) 0.28 (8K)
qwen72b 0.23 (8K) 0.29 (4K) 0.29 (4K) 0.29 (4K)
We evaluate the cheap-sampling estimator from Sec. 4 on tasks where model noise dominates: QA-IB and Summarization. For each candidate chunk size, we evaluate D&C on only $m\in\{3,5,10\}$ randomly selected documents from the 128K-token setting and pick the best-performing chunk size to deploy. Tables 1 compares this low-budget selection against the exhaustive grid search over all documents and all chunk sizes. We observe that even with three to five samples per configuration, the selected chunk sizes are near-optimal and often exactly match the exhaustive-search optimum, delivering the same final scores while avoiding a full grid search. This trend aligns with our framework: superlinear and near-monotone model noise in $L$ drives a clear optimal region in chunk size; a handful of samples per configuration is enough to trace the coarse error contour and locate the optimum, yielding significant computational savings without sacrificing peak performance.